# Bidirectional predictive coding

### v1

@@ 1 Introduction
Visual inference plays a critical role in the brain, providing information processing for interpreting and interacting with the environment. Two frameworks have emerged to explain how the brain could implement visual inference. The first describes vision as a bottom-up discriminative process, which filters sensory inputs to extract behaviorally relevant information, akin to inference in feedforward neural networks [1]. The other formalises vision as a generative process where the brain generates top-down predictions of the sensory inputs, and develops neural responses as a result of Bayesian inference [2]. Experimental evidence suggests that visual perception may be a combination of both frameworks [3, 4]. However, most computational models for visual inference are either discriminative or generative. Models that are a combination of both, often have to compromise on either supervised or unsupervised learning performance, both of which are hallmarks of visual learning. Hence, it is unknown how the brain is able to perform these diverse tasks accurately.
In this work, we propose bidirectional predictive coding (bPC), a novel model of biological vision that can perform both generative and discriminative inference based on predictive coding. Traditional predictive coding can be formulated as either generative or discriminative, and in both cases it matches biological data and the performance of analogous machine learning models. It also employs Hebbian rules to learn the model parameters supporting either generative or discriminative inference, making it biologically more plausible than models relying on backpropagation (BP) to learn. In this work, we focus on evaluating and understanding bPC’s performance in a range of computational tasks, where predictive coding models with either generative or discriminative inference have been individually successful. Our key contributions can be summarised as follows:
@@
  • We show that bPC performs as well as its purely discriminative or generative counterparts on both supervised classification tasks and unsupervised representation learning tasks, and outperforms precedent hybrid predictive coding models;
  • We provide an explanation for the superior performance of bPC in both tasks, by comparing it to unidirectional models and showing that it learns an energy landscape which benefits detecting out-of-distribution data points;
  • We further show that bPC outperforms other PC models in two biologically relevant tasks, including multimodal learning and inference with occluded visual scenes, indicating its potential as a more faithful model for visual inference in the brain.
@@ 2 Background and related work
Figure 1: Model architectures used in our experiments. Arrows indicate the direction of prediction. Dashed arrows represent initialisation. Discriminative models (red) perform bottom-up predictions, generative models (blue) perform top-down predictions, and bidirectional models incorporate both. Models in the top row implement local predictive coding computations, whereas those in the bottom row rely on non-local, BP-based computations. Each prediction arrow corresponds to bidirectional connections in a neural circuit: one for prediction, the other for error backpropagation.
@@
Recent studies have focused on PC models with bottom-up predictions from sensory inputs to latent activities [5, 6], similar to feedforward neural networks. The energy function of this type of PC models is given by:
$E_{disc}(x,V)=\sum\nolimits_{l=2}^{L}\frac{1}{2}\|x_{l}-V_{l-1}f(x_{l-1})\|_{2}^{2}$ (1)
In this energy function, $x_{l}$ is the neural activity of layer $l$ where $x_{1}$ is set to the sensory data and $x_{L}$ is set to a label. The parameter $V_{l}$ represents the bottom-up weights of layer $l$ and $f$ is an activation function. We refer to this model as discriminative predictive coding (discPC) and illustrate it in the top panel of Figure 1A. In discPC, inference is performed by a forward pass from $x_{1}$ to $x_{l}$’s, and the learning rule of this model approximates BP, using only local computations and Hebbian plasticity [5, 7]. Recent work showed that discPC performs comparably to BP in classifying MNIST, Fashion-MNIST and CIFAR-10 [8] and outperforms BP at learning scenarios encountered by the brain, such as online and continual learning [6]. However, discPC lacks unsupervised learning capabilities due to the non-uniqueness of the solution to its generative dynamics [9].
@@
Classically, predictive coding (PC) uses top-down predictions from neural activities to sensory data, based on the hypothesis that the brain learns by minimizing the error between its predicted sensory inputs and the actual sensory input [10, 11]. With neurons arranged hierarchically, the objective function, or energy [6], of PC can be written as:
$E_{gen}(x,W)=\sum\nolimits_{l=1}^{L-1}\frac{1}{2}\|x_{l}-W_{l+1}f(x_{l+1})\|_{2}^{2}$ (2)
where $x_{1}$ is set to sensory data and $x_{L}$ can be clamped to a label for supervised learning or remain free for unsupervised learning. The parameters $W_{l}$ are top-down weights. We refer to this formulation of predictive coding as generative predictive coding (genPC). In genPC, inference of latent activities $x_{l}$ is performed by clamping $x_{1}$ to sensory inputs and iteratively updating $x_{l}$’s via gradient descent on the energy $E_{gen}$. genPC’s top-down predictions are illustrated in the top-left panel of Figure 1B.
@@
GenPC has explained various visual phenomena, such as extra-classical receptive field effects and repetition suppression [10, 12, 13, 14]. More recently, genPC has been employed to model associative memory [15] and unsupervised image generation [16, 17, 8]. The computational framework of genPC can be implemented within neural networks that utilize local computations and Hebbian plasticity [18]. However, genPC has poor performance in supervised learning tasks, where $x_{1}$ is clamped to the input image and the corresponding label is iteratively inferred in $x_{L}$ [19].
Hybrid predictive coding.
In this work, we benchmark our bPC model primarily against hybrid predictive coding (hybridPC). hybridPC is the only plausible PC model that incorporates both slow, iterative inference and fast, feedforward inference, to date. In HybridPC, fast inference is enabled by a bottom-up network added to the genPC network architecture, which serves only to initialize neural activities and does not affect their dynamics, as illustrated in the top-right panel of Figure 1B. The neural dynamics and weight updates of hybridPC minimise the following energy function:
where sg($\cdot$) is the stop gradient operation. HybridPC can perform both supervised and unsupervised learning. However, the supervised learning performance of hybridPC is much poorer than that of discPC. This is due to the stop gradient operation in the bottom-up stream of this model that prevents it from learning a discriminative model that maps well from $x_{1}$ to $x_{L}$ [19].
Other predictive coding models.
@@
Other formulations of PC models combine discriminative and generative characteristics. Sun and Orchard [9] noted the energy-minimising nature of PC models could theoretically allow image generation in discPC (Eq. 1), by clamping $x_{L}$ to a label and iteratively updating $x_{1}$. However, the generated images appear nonsensical due to the non-unique inferential dynamics solutions. Even with decay mechanisms to make the solution unique, the output generations remain highly noisy [9]. Noting that any neuron in a PC model can be updated to minimise an energy function regardless of the model architecture, Salvatori et al. [20] generalized the idea of Sun and Orchard [9] to a PC model where all neurons are interconnected with each other. However, it only slightly outperforms the classification performance of a linear classifier. Both studies suggest that dedicated circuits for discriminative and generative tasks are needed for satisfactory performance. Finally, Qiu et al. [21] proposed a PC model that performs feedforward, bottom-up inference using the transpose of the top-down weights. Although this model performs well in both supervised and unsupervised tasks, it is unlikely that the brain shares synaptic connections between distinct streams of processing, making it an implausible model compared to bPC.
@@
Computational models of cortical processing can be roughly categorized into discriminative, such as those that use a feedforward neural network to filter visual inputs [22, 23], and generative, which are rooted in the Bayesian Brain hypothesis [24, 25] and instantiated by theories like predictive coding [10, 11]. However, increasing evidence suggests that the brain takes a combined approach [26, 3, 4]. For example, the brain might use a discriminative approach to rapidly extract important features from visual scenes, and perform Bayesian inference on these features to form high-level perceptions such as object categories [27]. Only a small number of models explicitly integrates these two approaches. The Helmholtz machine, an early model trained with the wake-sleep algorithm [28], is analogous to a discriminative-generative process and can be seen as a precursor to variational autoencoders [29] trained via BP. The Symmetric Predictive Estimator [30] is a recent model designed to capture bidirectional processing in the brain. However, its effectiveness beyond simple toy tasks remains unclear.
3 Model: bidirectional predictive coding
@@
In contrast to genPC and discPC, bPC neurons perform both top-down and bottom-up predictions, as shown in the top panel of Figure 1C. Using the notations previously introduced, the energy function of bPC is given by:
$E(x,W,V)=\sum\nolimits_{l=1}^{L-1}\frac{\alpha_{gen}}{2}\|x_{l}-W_{l+1}f(x_{l+1})\|_{2}^{2}+\sum\nolimits_{l=2}^{L}\frac{\alpha_{disc}}{2}\|x_{l}-V_{l-1}f(x_{l-1})\|_{2}^{2}$ (4)
@@
where $W_{l}$ are the top-down weights, $V_{l}$ are the bottom-up weights, and $f$ is an activation function. $\alpha_{gen}$ and $\alpha_{disc}$ are scalar weighting constants, which are needed to account for magnitude differences in the errors of bottom-up and top-down predictions. The weighting constants can be viewed as learnable precision parameters [11]; however, they are tuned and kept constant in our implementation for simplicity. Unlike in hybridPC, no stop gradient is applied. In each trial of learning, the neural activities are first updated to minimise $E$ via gradient descent (neural dynamics) following:
@@
$\frac{dx_{l}}{dt}=-\nabla_{x}E=-\epsilon_{l}^{gen}-\epsilon_{l}^{disc}+f^{\prime}(x_{l})\odot\Big{(}W_{l}^{\top}\epsilon_{l-1}^{gen}+V^{\top}_{l}\epsilon_{l+1}^{disc}\Big{)},$ (5)
@@
denote the top-down and bottom-up prediction errors of neurons in layer $l$ respectively. $f^{\prime}$ denotes the derivative of the function $f$, and $\odot$ is the element-wise product.
After updating neural activities, the weights are updated to minimise $E$ via a gradient descent step:
$\Delta W_{l}\propto-\nabla_{W_{l}}E=\epsilon_{l-1}^{gen}f(x_{l})^{\top},\quad\Delta V_{l}\propto-\nabla_{V_{l}}E=\epsilon_{l+1}^{disc}f(x_{l})^{\top}$ (7)
Similar to discPC [5] and hybridPC [19], we initialise the model layers using a feedforward sweep from layer $x_{1}$ to $x_{L}$ along the bottom-up predictions. For instance, the second and third layers are initialised as $x_{2}=V_{1}f(x_{1})$ and $x_{3}=V_{2}f(V_{1}f(x_{1}))$ respectively. This initialisation strategy can be interpreted as a mechanism for fast amortized inference when a sensory input is initially encountered, which is also observed in the brain [31, 26].
@@ 3.1 Neural implementation
Figure 2: Neural implementation of bPC (left) and discPC (right).
The computations described in Eqs 5 and 7 can be implemented in a neural network with fully local computation and plasticity, as illustrated in Figure 2. This network contains value neurons, which encode $x_{l}$, error neurons, which represent prediction errors, and synaptic connections, which encode the model parameters. All computations are local, relying solely on pre- and post- synaptic activity. Value neuron dynamics depend on local error signals, their own activity, and incoming synaptic weights. Similarly, error neuron activity depends only on adjacent value neurons and synaptic weights. The plasticity of the weights is also Hebbian, as it is a product of pre- and post-synaptic activity. The local implementation of bPC inherits that of genPC or discPC (shown in Figure 2 right), although it has two distinct error neurons per value neuron for bottom-up and top-down prediction errors.

### v2

@@ 1 Introduction
Visual inference plays a critical role in the brain, providing information processing for interpreting and interacting with the environment. Two frameworks have emerged to explain how the brain could implement visual inference. The first describes vision as a bottom-up discriminative process, where sensory stimuli are progressively filtered through layered neural architectures, predicting behaviourally relevant outputs (Hubel & Wiesel, 1962). This framework resembles inference in feedforward neural networks commonly employed in machine learning for image classification. The other framework formalises vision as a generative process, where the brain constructs a probabilistic model of sensory inputs (Knill & Pouget, 2004). From this perspective, the brain learns a top-down generative model with priors over incoming sensory activity, and neural responses arise from the Bayesian inversion of this model, estimating posterior probabilities of brain states given sensory information. Various neural implementations of this inversion have been proposed, including variational methods (Friston, 2005; 2010) or sampling approaches (Fiser et al., 2010; Orbán et al., 2016; Haefner et al., 2016). Experimental evidence suggests that visual perception may be a combination of both frameworks (Teufel & Fletcher, 2020; Peters et al., 2024). Discriminative models explain the rapid initial responses to visual stimuli through efficient bottom-up processing of the brain (Peters et al., 2024), whereas generative models capture the probabilistic computations critical for optimally integrating noisy sensory inputs with prior knowledge, as displayed in perception and behaviour (Ernst & Banks, 2002; Wolpert et al., 1995; Knill & Richards, 1996; van Beers et al., 1999).
One computational model capable of capturing both generative and discriminative processing modes in the brain is predictive coding (PC). In its generative formulation, PC accounts for a broad range of biological data from the visual system (Srinivasan et al., 1982; Rao & Ballard, 1999; Hosoya et al., 2005) and is successful in tasks such as associative memory (Salvatori et al., 2021) and image generation (Oliviers et al., 2024). In its discriminative formulation, PC matches backpropagation-trained neural networks in image classification (Whittington & Bogacz, 2017; Pinchetti et al., 2025b) while also explaining neural data (Song et al., 2024). Crucially, PC relies on local computations and Hebbian learning rules consistent with biological constraints (Hebb, 1949; Posner et al., 1988; Lisman, 2017) and implementable in realistic neural circuits (Bogacz, 2017), making it a strong candidate for modelling both generative and discriminative learning in the brain. However, current PC formulations remain restricted to a single inference mode. Hybrid approaches that combine generative and discriminative versions exist (Tscshantz et al., 2023) but sacrifice performance in at least one domain, leaving open how a single PC model could flexibly support both inference modes as required for biological vision.
In this work, we propose bidirectional predictive coding (bPC), a novel model of biological vision. bPC provides a biologically grounded neural mechanism that explains how the brain can simultaneously perform generative and discriminative inference based on PC. We focus on evaluating and understanding bPC’s performance in a range of computational tasks, where PC models with either generative or discriminative inference have been individually successful. Our key contributions can be summarised as follows:
@@
  • We show that bPC performs as well as its purely discriminative or generative counterparts on both supervised classification tasks and unsupervised representation learning tasks, and outperforms precedent hybrid models;
  • We provide an explanation for the superior performance of bPC in both tasks, by showing that it learns an energy landscape that better captures the training data distribution than its unidirectional counterparts;
  • We further show that bPC outperforms other PC models in two biologically relevant tasks, including learning in a bimodal model architecture and inference with occluded visual scenes, indicating its potential as a more faithful model for visual inference in the brain.
@@ 2 Background and related work
Recent research has explored PC models that employ bottom-up predictions from sensory inputs to latent neural activities (Whittington & Bogacz, 2017; Song et al., 2024), analogous to traditional feedforward neural networks. These models are structured as hierarchical Gaussian models consisting of $L$ layers, each characterized by neural activity $x_{l}$, where $x_{1}$ corresponds to the sensory input and $x_{L}$ represents a label. PC learns this hierarchical Gaussian structure by minimizing an energy function corresponding to the negative joint log-likelihood of the model:
$E_{disc}(x,V)=\sum\nolimits_{l=2}^{L}\frac{1}{2}\|x_{l}-V_{l-1}f(x_{l-1})\|_{2}^{2},$ (1)
where $x$ denotes the set of neural states from $x_{1}$ to $x_{L}$, the parameters $V$ include the bottom-up weights $V_{l}$ of each layer $l$, and $f$ represents an activation function. We refer to this model as discriminative predictive coding (discPC) and illustrate it in the top panel of Figure 1A. In discPC, inference is performed by a forward pass from $x_{1}$ to $x_{l}$’s, and the learning rule of this model approximates BP, using only local computations and Hebbian plasticity (Whittington & Bogacz, 2017). Recent work showed that discPC performs comparably to BP in classifying MNIST, Fashion-MNIST and CIFAR-10 (Pinchetti et al., 2025b) and outperforms BP at learning scenarios encountered by the brain, such as online and continual learning (Song et al., 2024). However, discPC lacks unsupervised learning capabilities due to the non-uniqueness of the solution to its generative dynamics (Sun & Orchard, 2020).
Figure 1: Model architectures used in our experiments. Arrows indicate the direction of prediction and error propagation. Dashed arrows represent initialisation. Discriminative models (red) are parametrised by bottom-up mappings from sensory inputs to brain states, generative models (blue) are parametrised by top-down mappings from brain states to sensory inputs, and bidirectional models incorporate both directions. Models in the top row employ local computations and error propagation that are considered biologically plausible, whereas those in the bottom row utilise non-local, backpropagation-based error computations that lack biological plausibility.
@@
Classically, PC uses top-down predictions from neural activities to sensory data, based on the hypothesis that the brain learns by minimizing the error between its predicted sensory inputs and the actual sensory input (Rao & Ballard, 1999; Friston, 2005). With neurons arranged hierarchically, the negative joint log-likelihood can be written as:
$E_{gen}(x,W)=\sum\nolimits_{l=1}^{L-1}\frac{1}{2}\|x_{l}-W_{l+1}f(x_{l+1})\|_{2}^{2},$ (2)
@@
where $x_{1}$ is set to sensory data and $x_{L}$ can be clamped to a label for supervised learning or remain free for unsupervised learning. The parameters $W_{l}$ are top-down weights. We refer to this formulation of predictive coding as generative predictive coding (genPC). In genPC, inference of latent activities $x_{l}$ is performed by clamping $x_{1}$ to sensory inputs and iteratively updating $x_{l}$’s via gradient descent on the energy $E_{gen}$. genPC’s top-down predictions are illustrated in the top-left panel of Figure 1B. genPC has explained various visual phenomena, such as extra-classical receptive field effects and repetition suppression (Rao & Ballard, 1999; Hohwy et al., 2008; Auksztulewicz & Friston, 2016). More recently, genPC has been employed to model associative memory (Salvatori et al., 2021) and unsupervised image generation (Oliviers et al., 2024; Zahid et al., 2024). The computational framework of genPC can be implemented within neural networks that utilize local computations and Hebbian plasticity (Bogacz, 2017). However, genPC has poor performance in supervised learning tasks (Tscshantz et al., 2023).
Predictive coding models with mixed inference modes.
@@
In this work, we benchmark our bPC model primarily against hybrid predictive coding (hybridPC) (Tscshantz et al., 2023). In hybridPC, iterative inference proceeds in the same way as in genPC. However, an additional bottom-up network is introduced to provide a feedforward initialisation of neural activities. This network only sets the initial states and does not influence the subsequent dynamics, as illustrated in the top-right panel of Figure 1B. Although hybridPC performs both supervised and unsupervised learning, its supervised performance falls short of discPC. In this work, we show that bPC performs supervised learning on par with discPC, and we provide an explanation of hybridPC’s inferior performance through the lens of bPC (see SM H). Sun & Orchard (2020) noted the energy-minimising nature of PC models could theoretically allow image generation in discPC (Eq. 1), by clamping $x_{L}$ to a label and iteratively updating $x_{1}$. However, the generated images appear nonsensical due to the non-unique inferential dynamics solutions. Salvatori et al. (2022) generalised the idea of Sun & Orchard (2020) to a PC model where all neurons are interconnected with each other. However, it only slightly outperforms the classification performance of a linear classifier. Finally, Qiu et al. (2023) proposed a bidirectional PC model in which the connections between separate layers share the same weights, e.g. the bottom-up weights from layer l1l-1 to $l$ equal the top-down weights from layer $l$ to $l+1$. It is unlikely that the brain shares synaptic connections between separate layers of processing.
@@
Cortical processing models are often divided into discriminative, which use feedforward networks to filter inputs (Yamins et al., 2014; Fukushima, 1980), and generative, grounded in the Bayesian Brain hypothesis (Helmholtz, 1866; Knill & Pouget, 2004) and theories such as predictive coding (Rao & Ballard, 1999; Friston, 2005) or adaptive resonance (Grossberg, 2012). Growing evidence suggests the brain combines these approaches (Lamme & Roelfsema, 2000; Teufel & Fletcher, 2020; Peters et al., 2024): for example, it might employ a discriminative approach to rapidly extract important features, and perform Bayesian inference on these features to form high-level perceptions such as object categories (Yildirim et al., 2024). Few models explicitly integrate both. The wake-sleep algorithm used to train the Helmholtz machine (Dayan et al., 1995) is analogous to a generative-discriminative process. The Symmetric Predictive Estimator (Xu et al., 2017) models bidirectional processing, though only on toy tasks. A related bidirectional model based on information maximization (Bozkurt et al., 2023) improves biological plausibility but requires two inference phases per update. Interactive activation models (Rumelhart & McClelland, 1982) also perform bidirectional inference, but do not incorporate a learning mechanism.
Discriminative and generative models in machine learning.
Research in machine learning has likewise sought to integrate discriminative and generative pathways. Encoder–decoder architectures such as U-Nets (Ronneberger et al., 2015) combine bottom-up and top-down processing via skip connections, while models including VAVAE (Yang & Wang, 2025) and VAR (Tian et al., 2024) unify both pathways within a shared latent representation. Other approaches, exemplified by PredNet (Lotter et al., 2016), take more explicit inspiration from cortical circuitry to implement bidirectional information flow. Despite their strong empirical performance, these models depend on non-local learning signals and therefore do not provide a biologically plausible account of bidirectional learning in the brain.
3 Bidirectional predictive coding
@@
In contrast to genPC and discPC, bPC neurons perform both top-down and bottom-up predictions, as shown in the top panel of Figure 1C. bPC achieves this bidirectional inference by unifying the energy functions of genPC and discPC into a single formulation, enabling both generative and discriminative prediction within the same circuit. Using the notations previously introduced, the energy function of bPC given by:
$E(x,W,V)=\sum\nolimits_{l=1}^{L-1}\frac{\alpha_{gen}}{2}\|x_{l}-W_{l+1}f(x_{l+1})\|_{2}^{2}+\sum\nolimits_{l=2}^{L}\frac{\alpha_{disc}}{2}\|x_{l}-V_{l-1}f(x_{l-1})\|_{2}^{2},$ (3)
@@
where $W_{l}$ are the top-down weights and $V_{l}$ are the bottom-up weights. $\alpha_{gen}$ and $\alpha_{disc}$ are scalar weighting constants, which are needed to account for magnitude differences in the errors of bottom-up and top-down predictions. The weighting constants can be viewed as learnable precision parameters (Friston, 2005); however, they are tuned and kept constant in our implementation for simplicity.
In each trial of learning, we first initialise the layers of neural activity using a feedforward sweep from layer $x_{1}$ to $x_{L}$ along the bottom-up predictions, similar to discPC (Whittington & Bogacz, 2017). For instance, the second and third layers are initialised as $x_{2}=V_{1}f(x_{1})$ and $x_{3}=V_{2}f(V_{1}f(x_{1}))$ respectively. This initialisation strategy can be interpreted as a mechanism for fast amortised inference when a sensory input is initially encountered, which is also observed in the brain (Thorpe et al., 1996; Lamme & Roelfsema, 2000). All bPC experiments used this activity initialisation scheme. After, we update the neural activities to minimise $E$ via several gradient descent steps (neural dynamics) following:
@@
$\frac{dx_{l}}{dt}\propto-\nabla_{x}E=-\epsilon_{l}^{gen}-\epsilon_{l}^{disc}+f^{\prime}(x_{l})\odot\Big(W_{l}^{\top}\epsilon_{l-1}^{gen}+V^{\top}_{l}\epsilon_{l+1}^{disc}\Big)+\mathcal{N}(0,\sigma^{2}I),$ (4)
@@
denote the top-down and bottom-up prediction errors of neurons in layer $l$ respectively. $f^{\prime}$ denotes the derivative of the function $f$, and $\odot$ is the element-wise product. The normally distributed noise $\mathcal{N}$ is zero-mean, temporally uncorrelated, and independent across neurons. By default, we set $\sigma^{2}=0$, yielding deterministic dynamics that converge to the maximum a posteriori estimate. Unless stated otherwise, all experiments in this paper use these inference dynamics. Setting $\sigma^{2}=1$ induces stochastic dynamics that sample from the model’s posterior (Oliviers et al., 2024). These dynamics enable learning the distributions of sensory inputs.
After updating neural activities, the weights are updated to minimise $E$ via a single gradient descent step:
$\Delta W_{l}\propto-\nabla_{W_{l}}E=\epsilon_{l-1}^{gen}f(x_{l})^{\top},\quad\Delta V_{l}\propto-\nabla_{V_{l}}E=\epsilon_{l+1}^{disc}f(x_{l})^{\top}.$ (6)
Figure 2: Neural implementation of bPC (left) and discPC (right).
@@ 3.1 Neural implementation
The computations described in Eqs 4 and 6 can be implemented in a neural network with fully local computation and plasticity, as illustrated in Figure 2. This network contains value neurons, which encode $x_{l}$, error neurons, which represent prediction errors, and synaptic connections, which encode the model parameters. All computations are local, relying solely on pre- and post- synaptic activity. Value neuron dynamics depend on local error signals, their own activity, and incoming synaptic weights. Similarly, error neuron activity depends only on adjacent value neurons and synaptic weights. The plasticity of the weights is also Hebbian, as it is a product of pre- and post-synaptic activity. The local implementation of bPC inherits that of genPC or discPC (shown in Figure 2 right), although it has two distinct error neurons per value neuron for bottom-up and top-down prediction errors.